Determine whether the following individual events are overlapping or​ non-overlapping. Then find the probability of the combined event. Getting a sum of either 5​,11​,or 12 on a roll of two dice

Choose the correct answer below​ and, if​ necessary, fill in the answer box to complete your choice.

​(Type an integer or a simplified​ fraction.)

A. The individual events are overlapping. The probability of the combined event is _________?

or

B. The individual events are​ non-overlapping. The probability of the combined event is __________?
Answer #1

The individual events are not overlapping because they are mutually exclusive.

P(Sum is 5, 11 or 12) = 4/36 + 2/36 + 1/36 = 7/36

Therefore,

Option A; 7/36